The random sample shown below was selected from a normal distribution.

3,6,8,3,8,8

Complete parts a and b.

a. Construct a 99% confidence interval for the population mean μ.

(Round to two decimal places as​ needed.)
Answer #1

Solution:

Given that,

x x2
3 9
6 36
8 64
3 9
8 64
8 64
\sumx = 36 \sumx2 = 246

a ) The sample mean is \bar x

Mean \bar x   = \sum x / n)

= ( 3+6+8+3+8+8 / 6 )

= 36 / 6

= 6

Mean \bar x   = 6

The sample standard is S

S = \sqrt{} ( \sum x2 ) - (( \sum x )2 / n ) / 1 -n )

= \sqrt{} ( 246 ( -(36 )2 / 6 ) / 5

   = \sqrt{} 30 / 5

=\sqrt{}6

= 2.45

The sample standard is S = 2.45

b ) \bar x = 6

s = 2.45

n = 6

Degrees of freedom = df = n - 1 = 6- 1 = 5

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t\alpha /2,df = t0.005,5 =4.031

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 4.031* (2.45/ \sqrt 6 )

= 4.03

Margin of error = 1.00

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

6 - 4.03< \mu < 6 + 4.03

1.97< \mu < 10.03

(1.97, 10.03)